OpenStudy (gudetamas):

PLEASE HELP FOR MY FINALS NEXT WEEK >.< Find the equation of quadratic in that has an axis of symmetry of x=3 and includes (0,22) and (4,6)

OpenStudy (sunnnystrong):

Recall that: general form of Parabola is: \[y = ax^2 + bx + c\] Parabola goes through points (0,22) and (4,6) so: \[22=a(0)^2+b(0)+c\] so: c= 22**

OpenStudy (irishboy123):

super! and draw [sketch] the thing.....or what you know about it